# Detectors for Safe and Reliable LLMs: Implementations, Uses, and Limitations

### v2

@@ 3.2.1. Reliability and Efficiency
Recently, there has been a rise in using LLMs to evaluate LLMs (Kim et al., 2023; Chiang and Lee, 2023; Zheng et al., 2023; Zhu et al., 2023). However, other works have surfaced limitations to this LLM-based evaluation approach, noting issues such as the effect of inherent world knowledge in larger LLMs, potential biases specific to the LLM being used (Shen et al., 2023b; Wang et al., 2023), and the general expense of using LLMs which may be prohibitive (Samsi et al., 2023).
@@
On the other hand, detectors provide an efficient and transparent alternative. Due to their compact size, they can be run easily - with many not even needing a GPU. On transparency, it is an open problem regarding how to document the vast amount of data used in training LLMs; engineers have even resorted to adversarial approaches to recover such information (Nasr et al., 2023). Comparatively, we know the specific data that is used in training any given detector, by construction.
3.2.2. Automated Benchmarking
@@
There is significant work around safety evaluation of LLMs (Sun et al., 2024) and there exist many different associated benchmarks (Baldini et al., 2022; Parrish et al., 2022; Akyürek et al., 2022; Smith et al., 2022; Selvam et al., 2023; Dhamala et al., 2021; Nangia et al., 2020; Nadeem et al., 2020; Nagireddy et al., 2024; Kour et al., 2023). For the benchmarks that induce open generations, it is an open and an extremely hard problem to evaluate these generated outputs at scale. Detectors provide us with an automated, efficient, and reference-free metric based solution. For two such safety benchmarks which were internally developed, Atta-Q (Kour et al., 2023) and SocialStigmaQA (Nagireddy et al., 2024), several detectors were used to quantify the proportion of harmful generations from LLMs on these benchmarks. We note that detectors can be used as reference-free metrics on any standard text generation benchmarks - in addition to just harm-specific benchmarks. Therefore, the harm dimensions that these detectors represent can be added as additional evaluation criteria for LLMs.
3.3. Other aspects of LLM governance
@@
LLM governance combines policy, practices, and tools to oversee LLM model development, deployment, and use. In the earlier sections, we described ways in which detectors can be used post-deployment, but detectors play multiple roles in the governance of LLMs throughout their life-cycle. For example, during model training or fine-tuning, detectors are used to remove undesirable training data and improve model quality (Ngo et al., 2021) by reducing hallucinations (Raunak et al., 2021; Nie et al., 2019), improving semantic correctness (Dušek et al., 2019) and removing bias (Nagireddy et al., 2024). Detectors are used in steering output generation (Welleck et al., 2022) and augmenting data sources by using an existing detection mechanism to generate realistic and similar text that result in the opposite class (Madaan et al., 2021; Robeer et al., 2021) aiding in deeper understanding of LLM functioning.
@@
As a potential capability for IBM’s commercial foundation model governance platform, detectors provide a way to ensure that models meet policies that specify minimum model behavior requirements. For example, an organization may require that an LLM does not generate toxic output prior to deployment. Detectors also provide a quantitative way to track model drift over time and enable policies to be set such that corrective action can be taken when a model starts to operate outside a desired and pre-specified norm. In instances where a model is procured or acquired from a vendor, we use detectors as an evaluation mechanism to understand the risks that the acquired model may pose (Piorkowski et al., 2023). In summary, detectors provide a means to measure model behavior and establish policies and practices based on or in reaction to those measures.
4. Inherent Challenges
@@
At their core, many of the detectors intend to label social harms manifested in language. Their implementation entails making a judgment in determining (i.e., detecting) whether or not a human behavior or attribute constitutes harm. Disciplines such as information science, science and technology studies, and anthropology have developed extensive literature showing the inherent challenges that a system of classification imposes, calling attention to the sometimes invisible forces and categories built into technological infrastructures (Bowker and Star, 1999). This literature attests that constructing a category automatically entails valorizing a point of view and silencing another (Bowker and Star, 1999). If this is true, what are the implications for our efforts building the detectors?
@@
In this section, our intention is to make explicit the choices made in the construction of the detectors and to reflect on the contested definitiveness of classifying human attributes and behavior. In particular, there are two critical moments that reveal the material force that categories have in arranging algorithmic-based work. First, when we as practitioners define what constitutes harm, we are forced to conceptualize and reach a consensus on which social constructs are harmful or biased toward an individual (and which are not). These decisions materialize during data annotation and the construction of a ground truth from which to evaluate. A subsequent critical moment is when users interact with the system via the platform and categories (harmful vs. not harmful) are rendered visible to them. It is only through these interactions that users can formally assess the appropriateness of the categories made by practitioners in a precedent stage and context otherwise unbeknownst to them. Within both moments, many issues emerge which make defining categories of harm (social and otherwise) and subsequently assessing these categories inherently difficult. We highlight three of these challenges and related assumptions below, while acknowledging that these are neither exhaustive nor mutually exclusive.
@@
Challenge 1 - Discrepancies between contexts: The relevance and level of difficulty associated with accurately understanding the context of data production for their later categorization are not new problems and can be best observed within the context of content moderation (Caplan, 2018; Gillespie, 2018). While the capabilities of algorithms to categorize and identify topics have improved in the last decade, there has been extensive research showing that it often requires more than flagging themes to determine whether a piece of online content (e.g., text, image, video) has violated the standards of platform companies (Caplan, 2018). In content moderation, context, intent, linguistic, and cultural cues all matter (Leetaru, 2019; Caplan, 2018). For moderators to accurately and reliably determine whether a piece of content is in violation of the platform guidelines, they need to assess it considering the context of creation, background information and intention of the individual who made it, as well as the social conditions in which it was made and subsequently seen (Caplan, 2018).
Challenge 2 - Lexical variations: Previous research has shown that lexical variation might change depending on the context. For example, research on content moderation in pro-eating disorder online communities has found that members of these communities adopt lexical variants to engage with other members and circumvent enforced content moderation (Chancellor et al., 2016). As in Challenge 1, for moderators to accurately and reliably determine violations to permissible content, they might need to engage with the communities using the language and/or affected by the language or associated harm to better understand linguistic and usage nuances.
@@
Challenge 3 - Data annotation is always subjective: Data annotation has been defined as a sense-making practice of labeling a given dataset to make it categorizable and machine-readable (Miceli et al., 2020; Wang et al., 2022). However, previous research has shown that annotation is not a straightforward task, with multiple and varied interpretations which could be attached to each data instance (Khan and Hanna, 2022; Miceli et al., 2020, 2022). Data annotation is not agnostic, and it is unfortunately a fixed practice, in the sense that we create fixed categories of data through our datasets. In the areas of content moderation for hate speech, this work depends heavily on the local understanding of annotators who supplied the training data for the detector (Khan and Hanna, 2022). In toxicity detection, it is well known that model results are linked to the annotator’s perception of what is or is not toxic (Davani et al., 2023a, b; Sap et al., 2022) and that different annotators tend to disagree on how to annotate toxicity (Welbl et al., 2021; Aroyo et al., 2023). For moderators to consistently detect content violations, they must create and establish meaning around what constitutes a violation in the first place (i.e., ‘the ground truth’), and since this assignment of meaning cannot be separated from the individual (Muller et al., 2021; Aroyo and Welty, 2015) nor their practices and constraints (Miceli et al., 2020, 2022; Zhang et al., 2020; Passi and Jackson, 2018; Alvarado Garcia et al., 2023), moderators might need to reflect upon, discuss, and document what guides their interpretation of the data at hand (Miceli et al., 2020) and the data transformations that occur to make the harms more legible or ‘readable’ in computational terms (Elish and danah boyd, 2018).
@@
Gaps and Assumptions: Without adequate resources, time, or expertise to thoroughly address these challenges at the scale in which they are imposed, moderators may be forced to make assumptions and decisions about content that is or has been thoroughly de-contextualized. These might be positivist or descriptive in nature (Páez et al., 2012), in that moderators might treat the text as something that can be definitively proven or falsifiable, which carries with it both assumptions about ‘how the world is’ or ‘how things are’ (D’Ignazio and Klein, 2020) and assumptions that others agree with this interpretation (that there is always a ground truth or a single right answer for each data point (Aroyo and Welty, 2015)). Other assumptions might be normative or prescriptive (Páez et al., 2012), in that moderators carry with them their own ideas, experiences, biases, and sociocultural expectations pertaining to ‘how the world should be’, which influences whether or not they consider a given text to be harmful and in turn, through filtering, impacts what downstream users see as harmful or not. Other times, moderators may be faced with content that lacks necessary specificity, forcing them to make decisions about harm where there is not enough information - this may create highly strict or highly lenient annotation or filtering practices, or may result in very specific errors during evaluation (Balagoplan et al., 2023). Finally, there are often also larger speculative questions pertaining to the overall outcomes of the annotated text, where moderators might not be privy to future contexts of their labels’ use or may have very little decisional capacity or power to control future applications or flagged content. Examples of each of these can be found in Figure 3, where panels A and B show two LLM-generated sentences that were manually annotated to create a synthetic training dataset for the stigma detector (Appendix LABEL:sec:social-stigma-syth). Examples of questions pertaining to these assumptions and gaps are highlighted in both texts. Given all the unknown context, it can be appreciated how difficult it might be to assess whether or not stigma is present in these sentences, a challenge which extends from the training data all the way to evaluating responses and detector ability. For example, in Figure 2, we see a model’s response to a prompt with an associated stigma detection score, but it may be difficult to evaluate or explain its detection abilities confidently, given the aforementioned gaps and assumptions.
@@
Figure 3. Examples of synthetic data with associated questions, gaps, and assumptions.
4.1. A closer look into the stigma detector
@@
This section is organized as follows: we first start with a definition of stigma and highlight its ties to the three aforementioned challenges to social harm detection and related assumptions. Given these, we suggest future directions for us to improve detectors and provide recommendations for assessing their responses.
4.1.1. What is stigma?
“Stigma is defined as a social construct based on perceptions of visible or invisible marks or traits that discredit or disvalue individuals” (Maestre, 2020; Goffman, 2022). Stigma is operationalized between people, only when a trait or condition is considered undesirable within a social group (CORRIGAN, 2014; JONES and CORRIGAN, 2014; Meisenbach, 2010; Bracke et al., 2019). Thus, the notion of stigma is a contentious term in the sense that its definition depends on the prevalent values of a specific social context. What is labeled as stigma in one context might not be in another. This is in line with the issues mentioned in Challenge 1 about understanding nuances between different intentions and contexts of use. Moreover, stigma is inherently about structural and social power dynamics, historical contingencies, and human interactions - that is, it always involves a person or group of people who exhibit a particular attribute and those people who observe that attribute and categorize it as problematic (Goffman, 2022). Not everyone will view this attribute as stigmatizing in a moment, nor will they label it as a stigma consistently across contexts, communities, or time. This echoes both the issues mentioned in Challenge 2 regarding lexical variation of terms related to actual, potential, or previous social harms, as well as issues of subjectivity mentioned in Challenge 3.
When translating these challenges into considerations for the development of a robust stigma detector, it suggests that in order to train a model to recognize stigma-related language, we need to spend time examining specific lexicons within affected communities (or even within ‘instigating’ parties) in order to understand how toxic and triggering language and associated behaviors manifest (which has implications for moderation use cases) (Chancellor et al., 2016). Additionally, certain vulnerable communities might talk about a stigma differently, meaning the lexical manifestation of what could signal stigma in a text might/will vary in unanticipated ways (which has implications for data distribution). Similarly, sometimes the avoidance or absence of certain ‘obvious’, ‘explicit’, or ‘expected’ reflections of stigma can also, paradoxically or strategically, signal the presence of stigma, harm, or social norms, depending on the context and lexicon (which has implications for evaluation).
@@ 4.1.1. What is stigma?
4.1.2. Recommendation 1: Revisiting Conceptualizations
@@
  1. (1)
    Conducting further empirical research to articulate and document which stigmas will be appropriate to consider for the contexts in which the designed technology will be deployed. Rather than being broad, we suggest scoping and specifying the focus (for an example see (Landau et al., 2023)).
  2. (2)
@@
    Examine how those categories of stigma have been portrayed within text datasets, as well as how definitions of stigma might change depending on the context of deployment/application. For an example on gender categorization see (Scheuerman et al., 2020).
  4. (4)
    During our observations, we revisited the extensive literature on content moderation and mental health (Addison et al., 2023; Steiger et al., 2021), and we believe that there are several parallels between the goals of developing the detectors and the insights of content moderation research (Gillespie, 2018; Caplan, 2018). We will draw from that literature to develop more contextualized solutions.
4.1.3. Recommendation 2: Ground Truth and Data Annotation
@@
Due to the subtleties and nuances involved in describing or identifying harm, methods and considerations for annotation become vitally important to detection and similar capabilities. While there has been extensive research on best practices for annotation including documentation practices (Bender and Friedman, 2018; Pushkarna et al., 2022), reflexivity (Nathan et al., 2023; Miceli et al., 2021), and description of data annotators (Gray and Suri, 2019), we provide a couple top-of-mind suggestions below:
  1. (1)
@@
    Have multiple annotators label the data and if possible, try to recruit or involve annotators with different cultural backgrounds and life experiences to encourage diverse ways of approaching the phenomenon we are trying to label (Arhin et al., 2021).
  2. (2)
    Provide explicit and detailed instructions on what to label, including examples and definitions of what counts as harm, and consult with subject matter experts and lived experience experts in the definitions and chosen labels where appropriate. Document this process so that it is repeatable and transparent, and share these definitions in papers and along with any shared data or associated fact sheets (Pushkarna et al., 2022; Bender and Friedman, 2018; Arnold et al., 2019).
  3. (3)
@@
    Have methods to capture and document disagreement between annotators (Davani et al., 2022). There are many possibilities for how to work with or think through dissensus or differing annotation (which we do not have the space to elaborate on) (Scheuerman et al., 2021), but it is important that these moments are not erased, hidden, or immediately smoothed over (Plank, 2022).
  4. (4)
    If possible, be open to changing definitions, labels, or methods based on annotator or expert feedback, documenting each iteration and rationale.
4.1.4. Recommendation 3: Addressing Contextual and Lexical Discrepancy
As mentioned previously, the challenges associated with detecting social harms often revolve around differences between the context of dataset creation and context of dataset use (including lack of robustness to OOD data), whereby how a term or word is used will change in meaning or significance. While this is not a fully solvable problem, actions and efforts that might improve the applicability and benefit of detectors for specific use cases include:
  1. (1)
    Spending time assessing and documenting background information that might be relevant for appropriate detector use.
  2. (2)
    Positionality affects our analytical capabilities and decisions (Attia and Edge, 2017; Berger, 2015). Therefore, we recommend documenting the intentions and positionality of the people creating the dataset, along with conditions and constraints pertaining to its creation.
  3. (3)
    Engaging with the communities that use these terms or are affected by these terms to better understand nuances and subtleties of language use, as well as get feedback for whether detecting these terms might be beneficial or detrimental to people.
4.2. Why is this important?

### v3

@@ 3.2 Evaluation
Recently, there has been a rise in using LLMs to evaluate LLMs (Kim et al. 2023; Chiang and Lee 2023; Zheng et al. 2023; Zhu, Wang, and Wang 2023). However, other works have surfaced limitations to this LLM-based evaluation approach, noting issues such as the effect of inherent world knowledge in larger LLMs, potential biases specific to the LLM being used (Shen et al. 2023a; Wang et al. 2023), and the general expense of using LLMs which may be prohibitive (Samsi et al. 2023).
@@
On the other hand, detectors provide an efficient and transparent alternative. Due to their compact size, they can be run easily - with many not needing a GPU. On transparency, it is an open problem for how to document the vast amount of data used in training LLMs; engineers have resorted to adversarial approaches to recover such information (Nasr et al. 2023). Comparatively, we know the specific data that is used in training any given detector, by construction.
Automated Benchmarking
@@
There is significant work around safety evaluation of LLMs (Sun et al. 2024) and there exist many different associated benchmarks (Baldini et al. 2022; Parrish et al. 2022; Akyürek et al. 2022; Smith et al. 2022; Selvam et al. 2023; Dhamala et al. 2021; Nangia et al. 2020; Nadeem, Bethke, and Reddy 2020; Nagireddy et al. 2024a; Kour et al. 2023). For the benchmarks that induce open generations, it is an open and an extremely hard problem to evaluate these generated outputs at scale. Detectors provide us with an automated, efficient, and reference-free metric based solution. For two such safety benchmarks which were internally developed, Atta-Q (Kour et al. 2023) and SocialStigmaQA (Nagireddy et al. 2024a), several detectors were used to quantify the proportion of harmful generations from LLMs on these benchmarks. We note that detectors can be used as reference-free metrics on any standard text generation benchmarks - in addition to just harm-specific benchmarks. Therefore, the harm dimensions that these detectors represent can be added as additional evaluation criteria for LLMs.
3.3 Other aspects of LLM governance
@@
LLM governance combines policy, practices, and tools to oversee LLM model development, deployment, and use. In the earlier sections, we described ways in which detectors can be used post-deployment, but detectors play multiple roles in the governance of LLMs throughout their life-cycle. For example, during model training or fine-tuning, detectors are used to remove undesirable training data and improve model quality (Ngo et al. 2021) by reducing hallucinations (Raunak, Menezes, and Junczys-Dowmunt 2021; Nie et al. 2019), improving semantic correctness (Dušek, Howcroft, and Rieser 2019) and removing bias (Nagireddy et al. 2024a). Detectors are used in steering output generation (Welleck et al. 2022) and augmenting data sources by using an existing detection mechanism to generate realistic and similar text that result in the opposite class (Madaan et al. 2021; Robeer, Bex, and Feelders 2021) aiding in deeper understanding of LLM functioning.
@@
As a potential capability for IBM’s commercial foundation model governance platform, detectors provide a way to ensure that models meet policies that specify minimum model behavior requirements. For example, an organization may require that an LLM does not generate toxic output prior to deployment. Detectors also provide a quantitative way to track model drift over time and enable policies to be set such that corrective action can be taken when a model starts to operate outside a pre-specified norm. In instances where a model is procured or acquired from a vendor, we use detectors as an evaluation mechanism to understand the risks that the acquired model may pose (Piorkowski, Hind, and Richards 2023). In summary, detectors provide a means to measure model behavior and establish policies and practices based on or in reaction to those measures.
4 Inherent Challenges
@@
At their core, many detectors intend to label social harms manifested in language. Their implementation entails making a judgment in determining (i.e., detecting) whether a human behavior or attribute constitutes harm. Disciplines such as information science, science and technology studies, and anthropology have developed extensive literature showing the inherent challenges that a system of classification imposes, calling attention to the sometimes invisible forces and categories built into technological infrastructures (Bowker and Star 1999). This literature attests that constructing a category automatically entails valorizing a point of view and silencing another (Bowker and Star 1999). If this is true, what are the implications for our efforts building the detectors?
@@
In this section, our intention is to make explicit the choices made in the construction of the detectors and to reflect on the contested definitiveness of classifying human attributes and behavior. In particular, there are two critical moments that reveal the material force that categories have in arranging algorithmic-based work. First, when we as practitioners define what constitutes harm, we are forced to conceptualize and reach a consensus on which social constructs are harmful or biased toward an individual (and which are not). These decisions materialize during data annotation and the construction of a ground truth from which to evaluate. A subsequent critical moment is when users interact with the system via the platform and categories (harmful vs. not harmful) are rendered visible to them. It is only through these interactions that users can formally assess the appropriateness of the categories made by practitioners in a precedent stage and context otherwise unbeknownst to them. Within both moments, many issues emerge which make defining categories of harm (social and otherwise) and subsequently assessing these categories inherently difficult. We highlight two of these challenges and related assumptions below, while acknowledging that these are neither exhaustive nor mutually exclusive.
@@
Challenge 1 - Discrepancies between contexts: The relevance and level of difficulty associated with accurately understanding the context of data production for their later categorization are not new problems and can be best observed within the context of content moderation (Caplan 2018; Gillespie 2018). While the capabilities of algorithms to categorize and identify topics have improved in the last decade, there has been extensive research showing that it often requires more than flagging themes to determine whether a piece of online content (e.g., text, image, video) has violated the standards of platform companies (Caplan 2018). In content moderation, context, intent, linguistic, and cultural cues all matter (Leetaru 2019; Caplan 2018). For moderators to accurately and reliably determine whether a piece of content is in violation of the platform guidelines, they need to assess it considering the context of creation, background information and intention of the individual who made it, as well as the social conditions in which it was made and subsequently seen (Caplan 2018).
@@
Challenge 2 - Data annotation is always subjective: Data annotation has been defined as a sense-making practice of labeling a given dataset to make it categorizable and machine-readable (Miceli, Schuessler, and Yang 2020; Wang, Prabhat, and Sambasivan 2022). However, previous research has shown that annotation is not a straightforward task, with multiple and varied interpretations which could be attached to each data instance (Khan and Hanna 2022; Miceli, Schuessler, and Yang 2020; Miceli et al. 2022). Data annotation is not agnostic, and it is unfortunately a fixed practice, in the sense that we create fixed categories of data through our datasets. In the areas of content moderation for hate speech, this work depends heavily on the local understanding of annotators who supplied the training data for the detector (Khan and Hanna 2022). In toxicity detection, it is well known that model results are linked to the annotator’s perception of what is or is not toxic (Davani et al. 2023b, a; Sap et al. 2022) and that different annotators tend to disagree on how to annotate toxicity (Welbl et al. 2021; Aroyo et al. 2023). For moderators to consistently detect content violations, they must create and establish meaning around what constitutes a violation in the first place (i.e., ‘the ground truth’), and since this assignment of meaning cannot be separated from the individual (Muller et al. 2021; Aroyo and Welty 2015) nor their practices and constraints (Miceli, Schuessler, and Yang 2020; Miceli et al. 2022; Zhang, Muller, and Wang 2020; Passi and Jackson 2018; Alvarado Garcia et al. 2023), moderators might need to reflect upon, discuss, and document what guides their interpretation of the data at hand (Miceli, Schuessler, and Yang 2020) and the data transformations that occur to make the harms more legible or ‘readable’ in computational terms (Elish and danah boyd 2018).
@@
Gaps and Assumptions: Without adequate resources, time, or expertise to thoroughly address these challenges at the scale in which they are imposed, moderators may be forced to make assumptions and decisions about content that is or has been thoroughly de-contextualized. These might be positivist or descriptive in nature (Páez, Scott, and Morency 2012), in that moderators might treat the text as something that can be definitively proven or falsifiable, which carries with it both assumptions about ‘how the world is’ or ‘how things are’ (D’Ignazio and Klein 2020) and assumptions that others agree with this interpretation (that there is always a ground truth or a single right answer for each data point (Aroyo and Welty 2015)). Other assumptions might be normative or prescriptive (Páez, Scott, and Morency 2012), in that moderators carry with them their own ideas, experiences, biases, and sociocultural expectations pertaining to ‘how the world should be’, which influences whether or not they consider a given text to be harmful and in turn, through filtering, impacts what downstream users see as harmful or not. Other times, moderators may be faced with content that lacks necessary specificity, forcing them to make decisions about harm where there is not enough information - this may create highly strict or highly lenient annotation or filtering practices, or may result in very specific errors during evaluation (Balagoplan et al. 2023). Finally, there are often also larger speculative questions pertaining to the overall outcomes of the annotated text, where moderators might not be privy to future contexts of their labels’ use or may have very little decisional capacity or power to control future applications or flagged content. Examples of each of these can be found in Figure 3, where panels A and B show two LLM-generated sentences that were manually annotated to create a synthetic training dataset for the stigma detector (Appendix C). Examples of questions pertaining to these assumptions and gaps are highlighted in both texts. Given all the unknown context, it can be appreciated how difficult it is to assess whether stigma is present in these sentences, a challenge which extends from the training data all the way to evaluating responses and detector ability. For example, in Figure 2, we see a model’s response to a prompt with an associated stigma detection score, but it may be difficult to evaluate or explain its detection abilities confidently, given the aforementioned gaps and assumptions.
@@
Figure 3: Examples of synthetic data with associated questions, gaps, and assumptions.
4.1 A closer look into the stigma detector
@@
This section is organized as follows: we first start with a definition of stigma and highlight its ties to the aforementioned challenges to social harm detection and related assumptions. Given these, we suggest future directions for us to improve detectors and provide recommendations for assessing their responses.
What is stigma?
“Stigma is defined as a social construct based on perceptions of visible or invisible marks or traits that discredit or disvalue individuals” (Maestre 2020; Goffman 2022). Stigma is operationalized between people, only when a trait or condition is considered undesirable within a social group (CORRIGAN 2014; JONES and CORRIGAN 2014; Meisenbach 2010; Bracke, Delaruelle, and Verhaeghe 2019). Thus, the notion of stigma is a contentious term in the sense that its definition depends on the prevalent values of a specific social context. What is labeled as stigma in one context might not be in another. This is in line with the issues mentioned in Challenge 1 about understanding nuances between different intentions and contexts of use. Moreover, stigma is inherently about structural and social power dynamics, historical contingencies, and human interactions - that is, it always involves a person or group of people who exhibit a particular attribute and those people who observe that attribute and categorize it as problematic (Goffman 2022). Not everyone will view this attribute as stigmatizing in a moment, nor will they label it as a stigma consistently across contexts, communities, or time. This echoes issues of subjectivity mentioned in Challenge 2.
When translating these challenges into considerations for the development of a robust stigma detector, it suggests that in order to train a model to recognize stigma-related language, we need to spend time examining specific lexicons within affected communities (or even within ‘instigating’ parties) in order to understand how toxic and triggering language and associated behaviors manifest (which has implications for moderation use cases) (Chancellor et al. 2016). Additionally, certain vulnerable communities might talk about a stigma differently, meaning the lexical manifestation of what could signal stigma in a text might/will vary in unanticipated ways (which has implications for data distribution). Similarly, sometimes the avoidance or absence of certain ‘obvious’, ‘explicit’, or ‘expected’ reflections of stigma can also, paradoxically or strategically, signal the presence of stigma, harm, or social norms, depending on the context and lexicon (which has implications for evaluation).
@@
Recommendation 1: Revisiting Conceptualizations
@@
  1. 1.
    Conducting further empirical research to articulate and document which stigmas will be appropriate to consider for the contexts in which the designed technology will be deployed. Rather than being broad, we suggest scoping and specifying the focus (for an example see (Landau et al. 2023)).
  2. 2.
@@
    Examining how those categories of stigma have been portrayed within text datasets, as well as how definitions of stigma might change depending on the context of deployment/application.
Recommendation 2: Ground Truth and Data Annotation
@@
Due to the subtleties and nuances involved in describing or identifying harm, methods and considerations for annotation become vitally important to detection and similar capabilities. While there has been extensive research on best practices for annotation including documentation practices (Bender and Friedman 2018; Pushkarna, Zaldivar, and Kjartansson 2022), reflexivity (Nathan et al. 2023; Miceli et al. 2021), and description of data annotators (Gray and Suri 2019), we provide a couple top-of-mind suggestions below:
  1. 1.
@@
    Have multiple annotators label the data and if possible, try to recruit or involve annotators with different cultural backgrounds and life experiences to encourage diverse ways of approaching the phenomenon we are trying to label (Arhin et al. 2021).
  2. 2.
@@
    Have methods to capture and document disagreement between annotators (Davani, Díaz, and Prabhakaran 2022). There are many possibilities for how to work with or think through dissensus or differing annotation (Scheuerman, Hanna, and Denton 2021), but it is important that these moments are not erased, hidden, or immediately smoothed over (Plank 2022).
4.2 Why is this important?